# Generalized and Unified Equivalences between Hardness and Pseudoentropy

### v1

@@ 1.1 Background on Hardness–Randomness Equivalences
In summary, the Multicalibration Theorem comes close to providing a unified picture of hardness versus randomness, but it has two significant deficiencies. First, it has a factor of $2^{\ell}$ in the exponent of the runtime of $P$, which is unfortunately necessary (as we prove in Corollary 1.13). Second, it is only stated and used in [1] for nonuniform complexity, whereas the prior proofs of Theorems 1.2, 1.3 had uniform-complexity analogues [32, 34].
@@ 1.2 Our Contributions
In this work, we address the aforementioned deficiencies of both Theorems 1.4 and 1.5, and thereby provide a more unified and clarified understanding of hardness versus randomness. Specifically, we provide a substantial strengthening of Theorem 1.4 with the following features in Theorem 1.10:
@@
  • We state and prove a uniform-complexity generalization of our theorem in Theorem 1.11. This result encompasses the uniform-complexity versions of Theorems 1.2 and 1.3 as shown in [31, 34] and extends Theorem 1.4 to the uniform-complexity setting as well (which, to our knowledge, has not been achieved in prior work). The uniform-complexity setting plays and important role in [31], which uses the uniform-complexity version of Theorem 1.3 to give refined constructions of pseudorandom generators from one-way functions.
  • To justify the blow-up in our theorem depending on the complexity of (approximately) computing the (sub)gradients $\nabla\varphi$, we prove that a doubly-exponential dependence on $\ell$ is unavoidable if we do not place any computational efficiency assumptions on $\varphi$, even when $\Phi$ only contains a single convex function $\varphi$ with bounded subgradients (Theorem 1.12). This result implies that the doubly-exponential dependence is necessary for achieving multicalibration (as well as the weaker notion of calibrated multiaccuracy) even when there is only a single, extremely simple distinguisher (Corollary 1.13).
A key insight in our work is that full-fledged multicalibration is overkill for proving Theorems 1.2 and 1.3. Instead, it suffices to use a combination of two substantially weaker notions, multiaccuracy and weight-restricted calibration. The latter notion (weight-restricted calibration) is the part that is tailored to the specific family $\Phi$, imposing one constraint per function $\varphi\in\Phi$, leading to a complexity blow-up that depends on the maximum complexity of (the gradients of) the functions in $\Phi$. See Section 1.5 for a more detailed discussion about the relationships among these notions from the algorithmic fairness literature (multicalibration, multiaccuracy, and weight-restricted calibration).
In independent works, Casacuberta, Gopalan, Kanade, and Reingold [2] and Dwork and Tankala [6] demonstrated the power of (variants of) a similar notion termed calibrated multiaccuracy (see Section 1.5 for more discussion), with Casacuberta et al. [2] giving applications in agnostic learning and to proving Impagliazzo’s Hardcore Lemma (which, as mentioned earlier, is equivalent to Theorem 1.5 for the case $\ell=1$) and Dwork and Tankala [6] showing that it can substitute for multicalibration in the results of [25], which characterize computational indistinguishability under repeated samples.
@@ 1.3 Notation
To describe our results and techniques more precisely, it will be convenient to switch from random-variable notation as used above to a functional notion. Let $L$ be a positive integer, which represents $2^{\ell}$ in the discussion above. Let $\Delta_{L}\subseteq[0,1]^{L}$ be the set of probability distributions over a label space $[L]:=\{1,\ldots,L\}$, where we represent discrete probability distributions by their probability mass functions. Then our random variable $(X,Y)$ can be represented by the probability distribution $\mu$ of $X$ and a function $g^{*}:\{0,1\}^{n}\rightarrow\Delta_{L}$, where $g^{*}(x)$ is the distribution of $Y|_{X=x}$. Similarly, we can represent another random variable $Z$ jointly distributed with $X$ by a simulator $s:\{0,1\}^{n}\rightarrow\Delta_{L}$.
We normally would describe a distinguisher $d$ between $(X,Y)$ and $(X,Z)$ as a randomized function $d:\{0,1\}^{n}\times[L]\rightarrow\{0,1\}$. Instead, it will be convenient to describe them as functions $f:\{0,1\}^{n}\rightarrow[-1,1]^{L}$, where $f(x)_{y}$ represents $1-2\mathbb{E}[d(x,y)]$. Then computational indistinguishability between $(X,Y)$ and $(X,Z)$ can be conveniently described as follows.
@@
[…]
@@
$\mathsf{D}_{\varphi}(g^{*}\|g)$ ::=:= $\mathbb{E}_{x\sim\mu}\left[\mathsf{D}_{\varphi}(g^{*}(x)\|g(x)\right]$
::=:= $\mathbb{E}_{x\sim\mu}\left[\varphi(g^{*}(x))-\varphi(g(x))-\langle g^{*}(x)-g(x),\nabla\varphi(g(x))\rangle\right]$
@@
Theorem 1.9 indeed implies Theorem 1.4 by the following correspondences:
@@
The assumption $\ell=O(\log n)$ in Theorem 1.4 gives $L=\mathsf{poly}(n)$. We also have $T_{\varphi},T_{\varphi}^{\prime}=\mathsf{poly}(n)$ from the assumptions in Theorem 1.4. These conditions ensure that $T^{\prime}=\mathsf{poly}(T,n,1/\varepsilon)$ in Theorem 1.9, so we can use (1) to prove that Item 2 (hardness) in Theorem 1.4 implies Item 1 (pseudoentropy). Similarly, the reverse implication follows from (2).
@@ 1.4.1 Non-uniform Setting
Let $T>0$ be a time bound and let $\varepsilon\in(0,1)$ be an error parameter. Let $\Phi$ be a family of convex functions $\varphi:\Delta_{L}\to\mathbb{R}$. Assume that given $v\in\Delta_{L}$, the subgradient $\nabla\varphi(v)$ is bounded in $[-1,1]^{L}$ and can be computed to $\ell_{\infty}$ accuracy $\varepsilon/4$ in nonuniform time $T_{\Phi}$, for every $\varphi\in\Phi$. Then there exists
@@
In Theorem 1.10, the single simulator $s$ satisfies (4) for all $\varphi\in\Phi$, and it simultaneously achieves indistinguishability ($s\in\mathsf{Ind}(g^{*};T,\varepsilon)$) and computational efficiency ($s\in\mathsf{Time}(T^{\prime})$). Therefore, (4) is stronger than the following guarantee akin to (1) in Theorem 1.9:
@@
Here, the entropy maximizer $s$ on the left (the simulator) and the divergence minimizer $g$ on the right (the approximator) can be different and can change based on the choice of $\varphi$. By contrast, in (4), the single function $s$ simultaneously plays the roles of the simulator and the approximator simultaneously for all $\varphi\in\Phi$. This gives a rigorous confirmation for the aforementioned informal intuition from the Multicalibration Theorem (Theorem 1.5) that there is a close correspondence between the best efficiently computable approximator $g$ for $g^{*}$ (with respect to any $D_{\varphi}$) and the highest-entropy simulator $s$ that is indistinguishable from $g^{*}$.
The increase in complexity at (3) in Theorem 1.10 is polynomial in $L=2^{\ell}$, a significant improvement over the doubly exponential dependence on $\ell$ in Theorem 1.5. We achieve this by focusing on a class $\Phi$ of functions $\varphi$ whose subgradients $\nabla\varphi$ are efficiently computable (up to small $\ell_{\infty}$ error). This efficiency assumption is naturally satisfied by common entropy notions such as the Shannon entropy ($\varphi(v)=\sum_{i}v_{i}\ln v_{i}$), the min entropy ($\varphi(v)=\|v\|_{\infty}$), and the collision probability ($\varphi(v)=\sum_{i}v_{i}^{2}$, see Section 1.6). One caveat is that Theorem 1.10 also requires the subgradient $\nabla\varphi(v)$ to be bounded, which is not satisfied by the Shannon entropy when $v$ is close to the boundary of the probability simplex $\Delta_{L}$. However, this turns out to be a minor issue that can be easily addressed by slightly perturbing $v$ away from the boundary of $\Delta_{L}$ (see Section 5).
In Theorem 1.12 (which we present shortly), we show that some efficiency assumption on the convex functions $\varphi$ is necessary to avoid an exponential dependence on $L$ (i.e. to avoid a doubly-exponential dependence on $\ell$).
In contrast to Theorem 1.9, our Theorem 1.10 does not require $\varphi$ to be differentiable or strictly convex. When $\varphi$ is not differentiable, its subgradient at some $v\in\Delta_{L}$ may not be unique, and Theorem 1.10 holds as long as we use any fixed choice of subgradient to define the Bregman divergence $\mathsf{D}_{\varphi}$ (assuming that the subgradient is bounded and can be computed up to small $\ell_{\infty}$ error in time $T$). Therefore, our Theorem 1.10 encompasses both Theorems 1.2 and 1.3.
@@
Combining the two directions (4) and (5) of Theorem 1.10 in a different order, we get that for every time bound $T>0$ and every $g^{*}:\{0,1\}^{n}\to\Delta_{L}$, there exist $T^{\prime}=O(T+T_{\Phi})$,
@@
This means that the single low-complexity function $s\in\mathsf{Time}(T^{\prime\prime})$ achieves comparable or better approximation error than every $g\in\mathsf{Time}(T)$ w.r.t. every divergence notion $\mathsf{D}_{\varphi}$ for $\varphi\in\Phi$. In the recent learning theory literature, a function $s$ satisfying this property is termed an omnipredictor [11]. Our Theorem 1.10 thus recovers a result of [9] showing a construction of omnipredictors based on calibrated multiaccuracy. (See Section 1.5 for more discussion.)
@@ 1.4.2 Uniform Setting
In Theorem 1.11 below, we prove a uniform version of Theorem 1.10 that encompasses the uniform versions of Theorems 1.2 and 1.3 in [31, 34] and extends Theorem 1.4 also to the uniform setting (which has not been explicitly achieved in prior work). In the uniform setting, instead of having a single function $g^{*}:\{0,1\}^{n}\to[L]$, we are interested in a family of functions $g^{*}_{n,L}:\{0,1\}^{n}\to[L]$ defined for varying choices of input length $n$ and label space size $L$. Theorem 1.10 only allows us to construct an efficiently computable function $s$ satisfying (4) separately for each choice of $(n,L)$, so it does not guarantee the existence of a single uniform algorithm (e.g. a Turing machine) that efficiently computes $s$ for all choices of $(n,L)$. Our Theorem 1.11 addresses this limitation.
To state Theorem 1.11, we use terminology similar to the uniform versions of Theorems 1.2 and 1.3 in [31, 34], where we formalize a uniform distinguisher as an efficient uniform oracle $\mathcal{A}$ that, when given sample access to $g^{*},g:\{0,1\}^{n}\to\Delta_{L}$ on a distribution $\mu$, [^3] This means that $\mathcal{A}$ can obtain random samples $(x,y,y^{*})\in\{0,1\}^{n}\times[L]\times[L]$, where we first draw $x$ from $\mu$ and then independently draw $y\sim g(x),y^{*}\sim g^{*}(x)$. aims to output a function $f:\{0,1\}^{n}\to[-1,1]^{L}$ that witnesses the violation of Definition 1.6:
@@
We say $g$ is $(\mathcal{A},\varepsilon)$-distinguishable from $g^{*}$ if (6) holds with high probability, and otherwise we say $g$ is $(\mathcal{A},\varepsilon)$-indistinguishable. We defer the formal definition to Definitions 3.6 and 3.8.
Our goal is to construct an $(\mathcal{A},\varepsilon)$-indistinguishable function $s$ satisfying (4) using an efficient uniform algorithm that takes $\mathcal{A}$ as a subroutine, and we still want $s$ to satisfy (4) simultaneously for all $\varphi\in\Phi$. To achieve this, we assume that the function class $\{\nabla\varphi\}_{\varphi\in\Phi}$ is weakly-agnostically learnable, which is trivially satisfied when $\Phi$ only contains a single $\varphi$ whose subgradient $\nabla\varphi$ can be efficiently approximated (as in Theorems 1.2 and 1.3). Roughly speaking, we assume that there is a weak-agnostic learner $\mathcal{B}$ that, when given i.i.d. data points $(v_{1},z_{1}),\ldots,(v_{m},z_{m})\in\Delta_{L}\times[-1,1]^{L}$ drawn from an unknown distribution $\mu^{\prime}$, can efficiently find $r:\Delta_{L}\to[-1,1]$ such that $\mathbb{E}_{(v,z)\sim\mu^{\prime}}\langle z,r(v)\rangle$ is comparable with or larger than $\sup_{\varphi\in\Phi}\mathbb{E}_{(v,z)\sim\mu^{\prime}}\langle z,\nabla\varphi(v)\rangle$. We defer the formal definition to Definition 4.1.
Theorem 1.11 (Uniform version of Theorem 1.10; informal statement of Theorem 4.3).
@@
Theorem 1.11 encompasses the uniform versions of Theorems 1.2 and 1.3 in [31, 34] by choosing the entropy notion $\mathsf{H}_{\varphi}$ accordingly. In the uniform versions of Theorems 1.2 and 1.3, we assume that every function $s^{\prime}$ with $\mathsf{H}_{\varphi}(s^{\prime})-\mathsf{H}_{\varphi}(g^{*})$ exceeding some threshold $\beta$ is $(\mathcal{A},\varepsilon)$-distinguishable from $g^{*}$. Under this assumption, since the function $s$ from Theorem 1.11 is $(\mathcal{A},\varepsilon)$-indistinguishable from $g^{*}$, it must hold that $\mathsf{H}_{\varphi}(s)-\mathsf{H}_{\varphi}(g^{*})\leq\beta$, which implies that $\mathsf{D}_{\varphi}(g^{*}\|s)\leq\beta+\varepsilon_{1}$. In this case, the algorithm $\mathcal{S}$ from Theorem 1.11 becomes an efficient uniform algorithm that approximates $g^{*}$ within low $\mathsf{D}_{\varphi}$ error, exactly as needed to prove the uniform versions of Theorems 1.2 and 1.3. We remark that the uniform versions of Theorems 1.2 and 1.3 in [31, 34] also have the reverse direction that can also be extended to general entropy notions $\mathsf{H}_{\varphi}$ to get a uniform version of the reverse direction (5) of Theorem 1.10. We omit this result as this generalization can be proved straightforwardly using the same proof idea as its non-uniform counterpart.

### v2

@@ 1.1 Background on Hardness–Randomness Equivalences
In summary, the Multicalibration Theorem comes close to providing a unified picture of hardness versus randomness, but it has two significant deficiencies. First, it has a factor of $2^{\ell}$ in the exponent of the runtime of $P$, which is unfortunately necessary (as we prove in Corollary˜1.13). Second, it is only stated and used in [1] for nonuniform complexity, whereas the prior proofs of Theorems 1.2, 1.3 had uniform-complexity analogues [32, 34].
@@ 1.2 Our Contributions
In this work, we address the aforementioned deficiencies of both Theorems˜1.4 and 1.5, and thereby provide a more unified and clarified understanding of hardness versus randomness. Specifically, we provide a substantial strengthening of Theorem 1.4 with the following features in Theorem˜1.10:
@@
  • We state and prove a uniform-complexity generalization of our theorem in Theorem˜1.11. This result encompasses the uniform-complexity versions of Theorems˜1.2 and 1.3 as shown in [31, 34] and extends Theorem˜1.4 to the uniform-complexity setting as well (which, to our knowledge, has not been achieved in prior work). The uniform-complexity setting plays and important role in [31], which uses the uniform-complexity version of Theorem˜1.3 to give refined constructions of pseudorandom generators from one-way functions.
  • To justify the blow-up in our theorem depending on the complexity of (approximately) computing the (sub)gradients $\nabla\varphi$, we prove that a doubly-exponential dependence on $\ell$ is unavoidable if we do not place any computational efficiency assumptions on $\varphi$, even when $\Phi$ only contains a single convex function $\varphi$ with bounded subgradients (Theorem˜1.12). This result implies that the doubly-exponential dependence is necessary for achieving multicalibration (as well as the weaker notion of calibrated multiaccuracy) even when there is only a single, extremely simple distinguisher (Corollary˜1.13).
A key insight in our work is that full-fledged multicalibration is overkill for proving Theorems 1.2 and 1.3. Instead, it suffices to use a combination of two substantially weaker notions, multiaccuracy and weight-restricted calibration. The latter notion (weight-restricted calibration) is the part that is tailored to the specific family $\Phi$, imposing one constraint per function $\varphi\in\Phi$, leading to a complexity blow-up that depends on the maximum complexity of (the gradients of) the functions in $\Phi$. See Section˜1.5 for a more detailed discussion about the relationships among these notions from the algorithmic fairness literature (multicalibration, multiaccuracy, and weight-restricted calibration).
In independent works, Casacuberta, Gopalan, Kanade, and Reingold [2] and Dwork and Tankala [6] demonstrated the power of (variants of) a similar notion termed calibrated multiaccuracy (see Section˜1.5 for more discussion), with Casacuberta et al. [2] giving applications in agnostic learning and to proving Impagliazzo’s Hardcore Lemma (which, as mentioned earlier, is equivalent to Theorem 1.5 for the case $\ell=1$) and Dwork and Tankala [6] showing that it can substitute for multicalibration in the results of [25], which characterize computational indistinguishability under repeated samples.
@@ 1.3 Notation
To describe our results and techniques more precisely, it will be convenient to switch from random-variable notation as used above to a functional notion. Let $L$ be a positive integer, which represents $2^{\ell}$ in the discussion above. Let $\Delta_{L}\subseteq[0,1]^{L}$ be the set of probability distributions over a label space $L]:=\{1,\ldots,L\}$, where we represent discrete probability distributions by their probability mass functions. Then our random variable $(X,Y)$ can be represented by the probability distribution $\mu$ of $X$ and a function $g^{*}:\{0,1\}^{n}\rightarrow\Delta_{L}$, where $g^{*}(x)$ is the distribution of $Y|_{X=x}$. Similarly, we can represent another random variable $Z$ jointly distributed with $X$ by a simulator $s:\{0,1\}^{n}\rightarrow\Delta_{L}$.
We normally would describe a distinguisher $d$ between $(X,Y)$ and $(X,Z)$ as a function $d:\{0,1\}^{n}\times[L]\rightarrow[0,1]$. Instead, it will be convenient to describe them as functions $f:\{0,1\}^{n}\rightarrow[-1,1]^{L}$, where the $y$-th coordinate of $f(x)$ (denoted by $f(x)_{y}$) represents 12d(x,y)1-2d(x,y). Then computational indistinguishability between $(X,Y)$ and $(X,Z)$ can be conveniently described as follows.
@@
where $\langle\cdot,\cdot\rangle$ denotes the standard inner product on $\mathbb{R}^{L}$. We write $\mathsf{Ind}_{\mu}(g^{*};T,\varepsilon)$ to denote the class of all functions $s:\{0,1\}^{n}\rightarrow\Delta_{L}$ that are $(T,\varepsilon)$-indistinguishable from $g^{*}$.
@@
$\mathsf{D}_{\varphi}(g^{*}\|g)$ $:=$ $\mathbb{E}_{x\sim\mu}\left[\mathsf{D}_{\varphi}(g^{*}(x)\|g(x)\right]$
$:=$ $\mathbb{E}_{x\sim\mu}\left[\varphi(g^{*}(x))-\varphi(g(x))-\langle g^{*}(x)-g(x),\nabla\varphi(g(x))\rangle\right]$
@@
Theorem˜1.9 indeed implies Theorem˜1.4 by the following correspondences:
@@
The assumption $\ell=O(\log n)$ in Theorem˜1.4 gives $L=\mathsf{poly}(n)$. We also have $T_{\varphi},T_{\varphi}^{\prime}=\mathsf{poly}(n)$ from the assumptions in Theorem˜1.4. These conditions ensure that $T^{\prime}=\mathsf{poly}(T,n,1/\varepsilon)$ in Theorem˜1.9, so we can use (1) to prove that Item˜2 (hardness) in Theorem˜1.4 implies Item˜1 (pseudoentropy). Similarly, the reverse implication follows from (2).
@@ 1.4.1 Non-uniform Setting
Let $T>0$ be a time bound and let $\varepsilon\in(0,1)$ be an error parameter. Let $\Phi$ be a family of convex functions $\varphi:\Delta_{L}\to\mathbb{R}$. Assume that given $v\in\Delta_{L}$, the subgradient $\nabla\varphi(v)$ is bounded in $1,1]^{L}$ and can be computed to $\ell_{\infty}$ accuracy $\varepsilon/4$ in nonuniform time $T_{\Phi}$, for every $\varphi\in\Phi$. Then there exists
@@
In Theorem˜1.10, the single simulator $s$ satisfies (4) for all $\varphi\in\Phi$, and it simultaneously achieves indistinguishability ($s\in\mathsf{Ind}(g^{*};T,\varepsilon)$) and computational efficiency ($s\in\mathsf{Time}(T^{\prime})$). Therefore, (4) is stronger than the following guarantee akin to (1) in Theorem˜1.9:
@@
Here, the entropy maximizer $s$ on the left (the simulator) and the divergence minimizer $g$ on the right (the approximator) can be different and can change based on the choice of $\varphi$. By contrast, in (4), the single function $s$ simultaneously plays the roles of the simulator and the approximator simultaneously for all $\varphi\in\Phi$. This gives a rigorous confirmation for the aforementioned informal intuition from the Multicalibration Theorem (Theorem˜1.5) that there is a close correspondence between the best efficiently computable approximator $g$ for $g^{*}$ (with respect to any $D_{\varphi}$) and the highest-entropy simulator $s$ that is indistinguishable from $g^{*}$.
The increase in complexity at (3) in Theorem˜1.10 is polynomial in $L=2^{\ell}$, a significant improvement over the doubly exponential dependence on $\ell$ in Theorem˜1.5. We achieve this by focusing on a class $\Phi$ of functions $\varphi$ whose subgradients $\nabla\varphi$ are efficiently computable (up to small $\ell_{\infty}$ error). This efficiency assumption is naturally satisfied by common entropy notions such as the Shannon entropy ($\varphi(v)=\sum_{i}v_{i}\ln v_{i}$), the min entropy ($\varphi(v)=\|v\|_{\infty}$), and the collision probability ($\varphi(v)=\sum_{i}v_{i}^{2}$, see Section˜1.6). One caveat is that Theorem˜1.10 also requires the subgradient $\nabla\varphi(v)$ to be bounded, which is not satisfied by the Shannon entropy when $v$ is close to the boundary of the probability simplex $\Delta_{L}$. However, this turns out to be a minor issue that can be easily addressed by slightly perturbing $v$ away from the boundary of $\Delta_{L}$ (see Section˜5).
In Theorem˜1.12 (which we present shortly), we show that some efficiency assumption on the convex functions $\varphi$ is necessary to avoid an exponential dependence on $L$ (i.e. to avoid a doubly-exponential dependence on $\ell$).
In contrast to Theorem˜1.9, our Theorem˜1.10 does not require $\varphi$ to be differentiable or strictly convex. When $\varphi$ is not differentiable, its subgradient at some $v\in\Delta_{L}$ may not be unique, and Theorem˜1.10 holds as long as we use any fixed choice of subgradient to define the Bregman divergence $\mathsf{D}_{\varphi}$ (assuming that the subgradient is bounded and can be computed up to small $\ell_{\infty}$ error in time $T$). Therefore, our Theorem˜1.10 encompasses both Theorems˜1.2 and 1.3.
@@
Combining the two directions (4) and (5) of Theorem˜1.10 in a different order, we get that for every time bound $T>0$ and every $g^{*}:\{0,1\}^{n}\to\Delta_{L}$, there exist $T^{\prime}=O(T+T_{\Phi})$,
@@
This means that the single low-complexity function $s\in\mathsf{Time}(T^{\prime\prime})$ achieves comparable or better approximation error than every $g\in\mathsf{Time}(T)$ w.r.t. every divergence notion $\mathsf{D}_{\varphi}$ for $\varphi\in\Phi$. In the recent learning theory literature, a function $s$ satisfying this property is termed an omnipredictor [11]. Our Theorem˜1.10 thus recovers a result of [9] showing a construction of omnipredictors based on calibrated multiaccuracy. (See Section˜1.5 for more discussion.)
@@ 1.4.2 Uniform Setting
In Theorem˜1.11 below, we prove a uniform version of Theorem˜1.10 that encompasses the uniform versions of Theorems˜1.2 and 1.3 in [31, 34] and extends Theorem˜1.4 also to the uniform setting (which has not been explicitly achieved in prior work). In the uniform setting, instead of having a single function $g^{*}:\{0,1\}^{n}\to[L]$, we are interested in a family of functions $g^{*}_{n,L}:\{0,1\}^{n}\to[L]$ defined for varying choices of input length $n$ and label space size $L$. Theorem˜1.10 only allows us to construct an efficiently computable function $s$ satisfying (4) separately for each choice of $(n,L)$, so it does not guarantee the existence of a single uniform algorithm (e.g. a Turing machine) that efficiently computes $s$ for all choices of $(n,L)$. Our Theorem˜1.11 addresses this limitation.
To state Theorem˜1.11, we use terminology similar to the uniform versions of Theorems˜1.2 and 1.3 in [31, 34], where we formalize a uniform distinguisher as an efficient uniform oracle $\mathcal{A}$ that, when given sample access to $g^{*},g:\{0,1\}^{n}\to\Delta_{L}$ on a distribution $\mu$, [^3] This means that $\mathcal{A}$ can obtain random samples $(x,y,y^{*})\in\{0,1\}^{n}\times[L]\times[L]$, where we first draw $x$ from $\mu$ and then independently draw $y\sim g(x),y^{*}\sim g^{*}(x)$. aims to output a function $f:\{0,1\}^{n}\to[-1,1]^{L}$ that witnesses the violation of Definition˜1.6:
@@
We say $g$ is $(\mathcal{A},\varepsilon)$-distinguishable from $g^{*}$ if (6) holds with high probability, and otherwise we say $g$ is $(\mathcal{A},\varepsilon)$-indistinguishable. We defer the formal definition to Definitions˜3.6 and 3.8.
Our goal is to construct an $(\mathcal{A},\varepsilon)$-indistinguishable function $s$ satisfying (4) using an efficient uniform algorithm that takes $\mathcal{A}$ as a subroutine, and we still want $s$ to satisfy (4) simultaneously for all $\varphi\in\Phi$. To achieve this, we assume that the function class $\{\nabla\varphi\}_{\varphi\in\Phi}$ is weakly-agnostically learnable, which is trivially satisfied when $\Phi$ only contains a single $\varphi$ whose subgradient $\nabla\varphi$ can be efficiently approximated (as in Theorems˜1.2 and 1.3). Roughly speaking, we assume that there is a weak-agnostic learner $\mathcal{B}$ that, when given i.i.d. data points $(v_{1},z_{1}),\ldots,(v_{m},z_{m})\in\Delta_{L}\times[-1,1]^{L}$ drawn from an unknown distribution $\mu^{\prime}$, can efficiently find $r:\Delta_{L}\to[-1,1]$ such that $\mathbb{E}_{(v,z)\sim\mu^{\prime}}\langle z,r(v)\rangle$ is comparable with or larger than $\sup_{\varphi\in\Phi}\mathbb{E}_{(v,z)\sim\mu^{\prime}}\langle z,\nabla\varphi(v)\rangle$. We defer the formal definition to Definition˜4.1.
Theorem 1.11 (Uniform version of Theorem˜1.10; informal statement of Theorem˜4.3).
@@
Theorem˜1.11 encompasses the uniform versions of Theorems˜1.2 and 1.3 in [31, 34] by choosing the entropy notion $\mathsf{H}_{\varphi}$ accordingly. In the uniform versions of Theorems˜1.2 and 1.3, we assume that every function $s^{\prime}$ with $\mathsf{H}_{\varphi}(s^{\prime})-\mathsf{H}_{\varphi}(g^{*})$ exceeding some threshold $\beta$ is $(\mathcal{A},\varepsilon)$-distinguishable from $g^{*}$. Under this assumption, since the function $s$ from Theorem˜1.11 is $(\mathcal{A},\varepsilon)$-indistinguishable from $g^{*}$, it must hold that $\mathsf{H}_{\varphi}(s)-\mathsf{H}_{\varphi}(g^{*})\leq\beta$, which implies that $\mathsf{D}_{\varphi}(g^{*}\|s)\leq\beta+\varepsilon_{1}$. In this case, the algorithm $\mathcal{S}$ from Theorem˜1.11 becomes an efficient uniform algorithm that approximates $g^{*}$ within low $\mathsf{D}_{\varphi}$ error, exactly as needed to prove the uniform versions of Theorems˜1.2 and 1.3. We remark that the uniform versions of Theorems˜1.2 and 1.3 in [31, 34] also have the reverse direction that can also be extended to general entropy notions $\mathsf{H}_{\varphi}$ to get a uniform version of the reverse direction (5) of Theorem˜1.10. We omit this result as this generalization can be proved straightforwardly using the same proof idea as its non-uniform counterpart.